# Weak–strong uniqueness for bi-fluid compressible system with algebraic closure

### v1

@@ 5 The weak–strong uniqueness principle
Let us now more precisely formulate our main theorem comparing global weak solutions on the time interval $(0,T)$ with the local strong solutions.
@@
be two distributional solutions to the transport equation (3.4). Suppose, in addition that
$\nabla\beta\in L^{\infty}((0,T)\times\Omega;\mathbb{R}^{3}),\quad\operatorname{div}{\bf v}\in L^{1}(0,T;L^{\infty}(\Omega)).$
Then, for any $\delta>0$ and for all $\tau\in[0,T]$, we have
@@
Proof.  To justify this lemma rigorously, one may use the DiPerna-Lions regularizing technique [6], see also Lemma 7.3 in [15] for the same argument applied to the pure transport equations for $\alpha,\beta$. In our case the equations of $\alpha,\beta$ contain additional nonlinear terms, so the analysis is more involved. We have
@@
Multiplying this equation by $(\alpha-\beta)$, we obtain
$\int_{\Omega}\frac{1}{2}(\alpha-\beta)^{2}\ {\rm d}{x}-\frac{1}{2}\int_{\Omega}\operatorname{div}{\bf v}(\alpha-\beta)^{2}\ {\rm d}{x}$
@@
We may choose $\delta>0$ arbitrarily small, such that
$\int_{\Omega}\frac{1}{2}(\alpha-\beta)^{2}\ {\rm d}{x}$
@@
$\quad+C_{\delta}\left(\|\alpha-\beta\|_{L^{\infty}(\Omega)}^{2}+\|\nabla\beta\|_{L^{\infty}(\Omega)}^{2}+\|\operatorname{div}{\bf v}\|^{2}_{L^{\infty}(\Omega)}\right)\int_{\Omega}(\alpha-\beta)^{2}\ {\rm d}{x},$
where we also used the crucial fact that $\omega_{\alpha}$ is uniformly bounded. The statement thus follows from integrating the above inequality with respect to $t\in(0,\tau)$. $\Box$
With Lemma 5.2 at hand, we may proceed to the proof of Theorem 5.1. We will prove it by estimating $J_{1},\ldots,J_{10}$ term by term.
@@
The second term might be estimated by $\int_{0}^{\tau}C(t)\bar{\cal E}(R,Q,{\bf u}|{\widetilde{R}},{\widetilde{Q}},{\bf v})\,{\rm d}t$ with $C$ depending on $\|\nabla{\bf v}\|_{L^{\infty}(\Omega)}$. To estimate the first term, we use
@@
and similarly for $Q-{\widetilde{Q}}$. Thus together we obtain
@@
where $C>0$ depends only on $\|({\widetilde{\varrho}_{+}},{\widetilde{\varrho}_{-}})\|_{L^{\infty}(\Omega)}$ and on $\|\alpha\|_{L^{\infty}(\Omega)}$. Therefore
@@
with $C_{\delta}$ depending additionally on $\|\partial_{t}{\bf v}+{\bf v}\cdot\nabla{\bf v}\|_{L^{\infty}(\Omega)}$. For the residual part of this term in $J_{1}$, considering the small and large values of ${\varrho_{-}},{\varrho_{+}}$ separately, we get
@@
with $C_{\delta}$ proportional to $\|\partial_{t}{\bf v}+{\bf v}\cdot\nabla{\bf v}\|_{L^{\infty}(\Omega)}^{2}$. Ultimately we have
@@
where $C_{\delta}$ depends on $\|\partial_{t}{\bf v}+{\bf v}\cdot\nabla{\bf v}\|_{L^{\infty}(\Omega)}$ and on $\|\nabla{\bf v}\|_{L^{\infty}(\Omega)}$.
Estimate of $J_{2}$ and $J_{3}$. Again, both the two terms will be considered on the essential and residual parts of the domain. Using Taylor expansion for $J_{2}$, we obtain
@@
where $C$ depends on $\|\alpha\|_{L^{\infty}(\Omega)}$ and $\|\operatorname{div}{\bf v}\|_{L^{\infty}(\Omega)}$. The estimate for the essential part of $J_{3}$ is similar. The residual part is bounded straight from (5). Summarizing, we have
@@
where $C_{\delta}$ depends on $\|{\widetilde{\varrho}_{+}}\|_{W^{1,\infty}(\Omega)}$. Moreover,
@@
with $C_{\delta}$ depending on $\|{\widetilde{\varrho}_{-}}\|_{W^{1,\infty}(\Omega)}$.
Estimate of $J_{6}$, $J_{7}$ and $J_{8}$. Here, all the terms vanish thanks to the algebraic closure. Indeed, we have
$J_{6}+J_{7}=$ $\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\left(\partial_{t}\alpha+{\bf u}\cdot\nabla\alpha\right){p_{+}}({\widetilde{\varrho}_{+}})\ {\rm d}{x}\,{\rm d}t$
$+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\Big(\partial_{t}(1-\alpha)+{\bf u}\cdot\nabla(1-\alpha)\Big){p_{-}}({\widetilde{\varrho}_{-}})\ {\rm d}{x}\,{\rm d}t$
$=$ $\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\left(\partial_{t}\alpha+{\bf u}\cdot\nabla\alpha\right){p_{+}}({\widetilde{\varrho}_{+}})\ {\rm d}{x}\,{\rm d}t$
$-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\Big(\partial_{t}\alpha+{\bf u}\cdot\nabla\alpha\Big){p_{-}}({\widetilde{\varrho}_{-}})\ {\rm d}{x}\,{\rm d}t$
$=$  $\,0,$
@@
$J_{8}=$ $-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}({\bf v}-{\bf u})\cdot\Big(\nabla\beta\,{p_{+}}({\widetilde{\varrho}_{+}})+\nabla(1-\beta)\,{p_{-}}({\widetilde{\varrho}_{-}})\Big)\ {\rm d}{x}\,{\rm d}t$
@@
Estimate of $J_{9}$ and $J_{10}$. To estimate the remaining terms, we couple them with terms
@@
$J_{9}$ $-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\alpha{p_{+}}({\widetilde{\varrho}_{+}})\frac{(\gamma-1)(1-\alpha)}{\gamma(1-\alpha)+\alpha}\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t$
@@
Notice that the last term $L_{2}$ can be already estimated, because the difference in the bracket is controlled by $C|\alpha-\beta|$, with $C>0$ depending on $\|(\alpha,\beta)\|_{L^{\infty}(\Omega)}$. We thus have
@@
with $C>0$ depending additionally on $\|{\widetilde{\varrho}_{+}}\|_{L^{\infty}(\Omega)}$, and $\|\operatorname{div}{\bf v}\|_{L^{\infty}(\Omega)}$.
We now transfer similarly
$J_{10}$ $+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\alpha{p_{-}}({\widetilde{\varrho}_{-}})\frac{(\gamma-1)(1-\alpha)}{\gamma(1-\alpha)+\alpha}\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t$
@@
with $C$ depending on $\|{\widetilde{\varrho}_{-}}\|_{L^{\infty}(\Omega)}$, and $\|\operatorname{div}{\bf v}\|_{L^{\infty}(\Omega)}$.
Finally let us look at the terms $L_{1}$ and ${\emph{\L }}_{1}$. Using again the equality of pressures, we can write
@@
for $C_{\delta}$ depending on $\|\partial_{t}{\bf v}+{\bf v}\cdot\nabla{\bf v}\|_{L^{\infty}(\Omega)}$, $\|\nabla{\bf v}\|_{L^{\infty}(\Omega)}$, $\|{\widetilde{\varrho}_{+}}\|_{W^{1,\infty}(\Omega)}$ and $\|{\widetilde{\varrho}_{-}}\|_{W^{1,\infty}(\Omega)}$, as well as on the upper bounds of $\alpha$ and $\beta$, which are equal to $1$.
Summing this estimate up with (5.2) and using the Grönwall argument, we finish the proof of Theorem 5.1. In particular, our main Theorem 2.1 follows. $\Box$
Acknowledgement. The work of Y.L. was supported by National Natural Science Foundation of China (12571228), Natural Science Foundation of Anhui Province (2408085MA018). The work of E.Z. was supported by the EPSRC Early Career Fellowship no. EP/V000586/1.

### v2

@@ 5 The weak–strong uniqueness principle
Let us now more precisely formulate our main theorem comparing global weak solutions in the time interval $(0,T)$ with the local strong solutions.
@@
be two distributional solutions to the transport equation (3.4). Suppose, in addition, that
$\nabla\beta\in L^{\infty}((0,T)\times\Omega;\mathbb{R}^{3}),\quad\operatorname{div}{\bf v}\in L^{1}(0,T;L^{\infty}(\Omega))$
and $0\leq\alpha,\beta\leq 1$ almost everywhere in $(0,T)\times\Omega$. Then, for any $\delta>0$ and for all $\tau\in[0,T]$, we have
@@
with $C_{\delta}(t)$ integrable over $(0,T)$.
Proof.  To rigorously justify this lemma, one may use the DiPerna-Lions regularizing technique [8], see also Lemma 7.3 in [20] for the same argument applied to the pure transport equations for $\alpha,\beta$. In our case, the equations of $\alpha,\beta$ contain additional nonlinear terms, so the analysis is more involved. We have
@@
Multiplying this equation by $(\alpha-\beta)$, we obtain ($\omega_{\alpha}=\frac{(\gamma-1)(1-\alpha)}{\gamma(1-\alpha)+\alpha}\alpha$, similarly $\omega_{\beta}$)
$\int_{\Omega}\frac{1}{2}\partial_{t}(\alpha-\beta)^{2}\ {\rm d}{x}-\frac{1}{2}\int_{\Omega}\operatorname{div}{\bf v}(\alpha-\beta)^{2}\ {\rm d}{x}$
@@
We may choose $\delta>0$ arbitrarily small so that
$\int_{\Omega}\frac{1}{2}\partial_{t}(\alpha-\beta)^{2}\ {\rm d}{x}$
@@
$\quad+C_{\delta}\left(\|\alpha-\beta\|_{L^{\infty}(\Omega)}^{2}+\|\nabla\beta\|_{L^{\infty}(\Omega)}^{2}+\|\operatorname{div}{\bf v}\|_{L^{\infty}(\Omega)}\right)\int_{\Omega}(\alpha-\beta)^{2}\ {\rm d}{x},$
where we also used the crucial fact that $\omega_{\alpha}$ is uniformly bounded. The statement thus follows by integrating the above inequality with respect to $t\in(0,\tau)$. $\Box$
With Lemma 5.2 at hand, we may proceed to the proof of Theorem 5.1. We will prove it by estimating $J_{1},\ldots,J_{8}$ term by term.
@@
The second term can be estimated by $\int_{0}^{\tau}C(t)\bar{\cal E}(R,Q,{\bf u}|{\widetilde{R}},{\widetilde{Q}},{\bf v})\,{\rm d}t$ with $C(t)$ depending on $\|\nabla{\bf v}\|_{L^{\infty}(\Omega)}(t)$. To estimate the first term, we use
@@
and similarly for $Q-{\widetilde{Q}}$. Thus, together, we obtain
@@
where $C=C(t)>0$ depends only on $\|({\widetilde{\varrho}_{+}},{\widetilde{\varrho}_{-}})\|_{L^{\infty}(\Omega)}(t)$ and on $\|\alpha\|_{L^{\infty}(\Omega)}(t)$. Therefore,
@@
with $C_{\delta}(t)$ proportional to $\|\partial_{t}{\bf v}+{\bf v}\cdot\nabla{\bf v}\|^{2}_{L^{\infty}(\Omega)}(t)$. For the residual part of this term in $J_{1}$, considering the small and large values of ${\varrho_{-}},{\varrho_{+}}$ separately, we get
@@
with $C_{\delta}(t)$ proportional to $\|\partial_{t}{\bf v}+{\bf v}\cdot\nabla{\bf v}\|_{L^{\infty}(\Omega)}(t)$. Ultimately, we have
@@
where $C_{\delta}(t)$ depends on $\|\partial_{t}{\bf v}+{\bf v}\cdot\nabla{\bf v}\|_{L^{\infty}(\Omega)}(t)$ and on $\|\nabla{\bf v}\|_{L^{\infty}(\Omega)}(t)$ so that $C_{\delta}(t)$ is integrable over $(0,T)$.
Estimate of $J_{2}$ and $J_{3}$. Again, both terms will be considered on the essential and residual parts of the domain. Using Taylor expansion for $J_{2}$, we obtain
@@
where $C(t)$ is proportional to $\|\alpha\|_{L^{\infty}(\Omega)}(t)\|\operatorname{div}{\bf v}\|_{L^{\infty}(\Omega)}(t)$. The estimate for the essential part of $J_{3}$ is similar. The residual part is bounded directly from (5) and $C(t)$ is again integrable. In summary, we have
@@
where $C_{\delta}(t)$ depends on $\|{\widetilde{\varrho}_{+}}\|_{W^{1,\infty}(\Omega)}(t)$. Moreover,
@@
with $C_{\delta}(t)$ depending on $\|{\widetilde{\varrho}_{-}}\|_{W^{1,\infty}(\Omega)}(t)$.
@@
$J_{6}=$ $-\int_{0}^{\tau}\!\!\int_{\Omega}({\bf v}-{\bf u})\cdot\Big(\nabla\beta\,{p_{+}}({\widetilde{\varrho}_{+}})+\nabla(1-\beta)\,{p_{-}}({\widetilde{\varrho}_{-}})\Big)\ {\rm d}{x}\,{\rm d}t$
@@
Estimate of $J_{7}$ and $J_{8}$. To estimate the remaining terms, we couple them with terms
@@
$J_{7}$ $-\int_{0}^{\tau}\!\!\int_{\Omega}\alpha{p_{+}}({\widetilde{\varrho}_{+}})\frac{(\gamma-1)(1-\alpha)}{\gamma(1-\alpha)+\alpha}\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t$
@@
Notice that the last term $L_{2}$ can already be estimated, because the difference in brackets is controlled by $C|\alpha-\beta|$, with $C>0$ depending on $\|(\alpha,\beta)\|_{L^{\infty}(\Omega)}$. Thus, we have
@@
with $C(t)>0$ depending additionally on $\|{\widetilde{\varrho}_{+}}\|_{L^{\infty}(\Omega)}(t)$, and $\|\operatorname{div}{\bf v}\|_{L^{\infty}(\Omega)}(t)$. Note that $C(t)$ is integrable over $(0,T)$.
We now transfer in a similar way
$J_{8}$ $+\int_{0}^{\tau}\!\!\int_{\Omega}\alpha{p_{-}}({\widetilde{\varrho}_{-}})\frac{(\gamma-1)(1-\alpha)}{\gamma(1-\alpha)+\alpha}\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t$
@@
with $C(t)$ depending on $\|{\widetilde{\varrho}_{-}}\|_{L^{\infty}(\Omega)}(t)$ and $\|\operatorname{div}{\bf v}\|_{L^{\infty}(\Omega)}(t)$ so that it is integrable.
Finally, let us look at the terms $L_{1}$ and ${\emph{\L }}_{1}$. Using again the equality of pressures, we can write
@@
for $C_{\delta}(t)$ depending on $\|\partial_{t}{\bf v}+{\bf v}\cdot\nabla{\bf v}\|_{L^{\infty}(\Omega)}(t)$, $\|\nabla{\bf v}\|_{L^{\infty}(\Omega)}(t)$, $\|{\widetilde{\varrho}_{+}}\|_{W^{1,\infty}(\Omega)}(t)$ and $\|{\widetilde{\varrho}_{-}}\|_{W^{1,\infty}(\Omega)}(t)$, as well as on the upper limits of $\alpha$ and $\beta$, which are equal to $1$, so that $C_{\delta}$ is integrable over $(0,T)$.
Summing up this estimate with (5.2) and using the Grönwall argument, we finish the proof of Theorem 5.1. In particular, our main Theorem 2.1 follows. $\Box$
Acknowledgement. The work of Y.L. was supported by National Natural Science Foundation of China (12571228), Natural Science Foundation of Anhui Province (2408085MA018). The work of M.L.-M. was supported by the Gutenberg Research College and by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation)–project number 233630050–TRR 146 and project number 525853336–SPP 2410 “Hyperbolic Balance Laws: Complexity, Scales and Randomness”. She is also grateful to the Mainz Institute of Multiscale Modelling for supporting her research. The work of M.P. was partially supported by the Czech Science Foundation (GAČR), project No. 25-16592S. The work of E.Z. was supported by the EPSRC Early Career Fellowship no. EP/V000586/1.